OpenStudy (anonymous):

Find the circumference and the area of a circle with diameter . Use the value 3.14 for (pi sign) , and do not round your answers. Be sure to include the correct units in

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (johnweldon1993):

\[\large \text{Area} = \pi r^2\] \[\large \text{Circumference} = 2 \pi r\] \[\large \text{Radius} = \frac{\text{Diameter}}{2}\]

OpenStudy (anonymous):

so....

OpenStudy (johnweldon1993):

so.... post a diameter...

OpenStudy (anonymous):

The diameter is 8

OpenStudy (johnweldon1993):

Alright...So now with all those formulas.. \[\large \text{Radius} = \frac{8}{2}\] \[\large \text{Radius} = 4\] Now that you know that...plug that...and the fact that pi 3.14 into the equation I have posted above \[\large \text{Circumference} = 2(3.14)(4) = ?\] \[\large \text{Area} = (3.14)(4)^2 = ?\]

OpenStudy (anonymous):

dont uderstand

OpenStudy (johnweldon1993):

What don't you understand? To solve for the circumference you would do 2 times 3.14 times 4 And to solve for the area you do 3.14 times 4 squared

OpenStudy (anonymous):

im stupid so... i still dont uderstand wat u r saying

OpenStudy (anonymous):

sorry